# A Textbook Remedy for Domain Shifts: Knowledge Priors for Medical Image Analysis

### v1

@@ Abstract
While deep networks have achieved broad success in analyzing natural images, when applied to medical scans, they often fail in unexcepted situations. We investigate this challenge and focus on model sensitivity to domain shifts, such as data sampled from different hospitals or data confounded by demographic variables such as sex, race, etc, in the context of chest X-rays and skin lesion images. A key finding we show empirically is that existing visual backbones lack an appropriate prior from the architecture for reliable generalization in these settings. Taking inspiration from medical training, we propose giving deep networks a prior grounded in explicit medical knowledge communicated in natural language. To this end, we introduce Knowledge-enhanced Bottlenecks (KnoBo), a class of concept bottleneck models that incorporates knowledge priors that constrain it to reason with clinically relevant factors found in medical textbooks or PubMed. KnoBo uses retrieval-augmented language models to design an appropriate concept space paired with an automatic training procedure for recognizing the concept. We evaluate different resources of knowledge and recognition architectures on a broad range of domain shifts across 20 datasets. In our comprehensive evaluation with two imaging modalities, KnoBo outperforms fine-tuned models on confounded datasets by 32.4 % on average. Finally, evaluations reveal that PubMed is a promising resource for making medical models less sensitive to domain shift, outperforming other resources on both diversity of information and final prediction performance.
@@ 2 Related Work
Knowledge Rich Multimodal Reasoning. Knowledge plays an important role in clinical diagnosis [6]. Some multimodal tasks [82, 53, 74] require models to use explicit outside knowledge to make correct predictions. Previous methods [52, 50, 31] retrieve documents for each example from the external knowledge base as context for models to generate the answer. Our work focuses on leveraging knowledge in medical image classification. Retrieval-Augmented Generation [48, 23] has been shown to be beneficial for knowledge-intensive tasks [38], including biomedicine [20, 84]. The retrieved medical documents are either used as context during inference [88] or data for pretraining [94]. In constrast, we treat documents as background knowledge for large language models to build concept bottlenecks. Instead of retrieving documents for every input, we build a global knowledge prior from a medical document corpus, which is shared across all examples.
@@ 4 Knowledge-enhanced Bottlenecks
In this section, we present Knowledge-enhanced Bottlenecks (KnoBo), a class of CBMs that incorporate knowledge priors that address the failures we identified in Section 3. Figure 3 presents an overview of our method, from left to right, we optimize three terms: (1) Structure Prior (Sec 4.2) induces bottleneck structures from medical corpus to incorporate human knowledge as concepts, (2) Bottleneck Predictor (Sec 4.3) projects input image onto bottleneck concepts and then feed concept predictions into the linear layer for label prediction, and (3) Parameter Prior (Sec 4.4) aligns the learned parameters with known associative information to further enhance priors.
@@ 4.1 Problem Formulation
$\log P\left(y,C,\theta|I\right)=\underbrace{\log P\left(y|I,C,\theta\right)}_{\text{bottleneck predictor}}+\underbrace{\log P\left(C\right)}_{\text{structure prior}}+\underbrace{\log P\left(\theta\right)}_{\text{parameter prior}}$ (1)
@@
Algorithm 1 Retrival Augmented Iterative Concept Bottleneck Generation
@@ 4.2 Structural Prior
Given a background corpus $\mathcal{B}$ that spans various documents, we aim to identify a bottleneck structure containing concepts beneficial for classifying labels $y\in\mathcal{Y}$. As outlined in Algorithm 1, we use the class names $\mathcal{Y}$ as initial queries to retrieve relevant documents $\mathcal{B^{\prime}}\subset\mathcal{B}$. Large Language Models (LLMs) are then prompted to generate concepts using these retrieved documents as context: $C^{\prime}=\text{LLM}(\mathcal{B^{\prime}})$. These newly generated concepts are added to the bottleneck and used as new queries to retrieve additional documents. This iterative process continues to expand the bottleneck until a predetermined number of concepts $N_{C}$ is reached. Such concept structures have high likeleehood under the language model, conditioned on the background corpus, and the language model probability serves as an implicit prior.
@@ 5.1 Datasets
We evaluate on two groups of datasets for each modality: (1) the confounded datasets, which aim to assess the robustness of models by creating splits with spurious correlations; (2) the unconfounded datasets are randomly split to measure the models’ performance in natural settings.
Confounded Datasets. As illustrated on the left of Figure 1, we formulate the confounded datasets as binary classification tasks, where each class is confounded with one factor. The confounding combinations are reversed for in-domain (train and validation) and out-of-domain (test) splits.
@@ 5.3 Implementation Details
KnoBo Details. We select GPT-4 (gpt-4-0613) as the underlying LLM for retrieval-augmented concept generation (Sec 4.2). For training concept grounding functions (Sec 4.3), we opt for Flan-T5-XXL [10] to annotate clinical reports for each concept, considering cost-efficiency. Unless otherwise specified, KnoBo uses bottlenecks constructed from PubMed, each containing 150 concepts. Figure 5 shows our prompt, and during concept generation, we apply several heuristic filters (Appendix  B.3).
@@ 6.1 Main Results
KnoBo is more robust to domain shifts. Table 1 shows the results on 10 confounded datasets of X-ray and skin lesions. Black-box models excel at in-domain (ID) data but drop significantly on out-of-domain (OOD) data, especially in datasets confounded by hospitals/resources (NIH-CheXpert and ISIC-hospital), which can be common when collecting medical datasets [12, 81]. KnoBo outperforms baselines in OOD and domain-average accuracy by large margins, ranking top-1 in eight datasets and second-best in the other two. End-to-end models (ViT-L/14, DenseNet) exhibit larger domain gaps than linear probes, as they have more parameters to optimize performance on in-domain data and capture spurious correlations. Shaping the visual representations with knowledge (LSL) improves robustness but underperforms KnoBo, with lower ID, OOD, and average performance across most datasets. PCBM-h combines interpretable and black-box predictions but exhibits behaviors similar to black-box models with severe drops across domains. Unlike KnoBo, which uses medical documents to create one global bottleneck for each modality, LaBo builds a bottleneck for each dataset using the in-domain data, which can be biased and affected by confounding factors, and so performs more poorly. In summary, KnoBo mitigates the catastrophic failures in domain shifts encountered by black-box models and is more robust against various confounding factors across modalities.
@@ 6.2 Analysis
In this section, we compare the bottlenecks constructed from different knowledge resources. We evaluate the impact of each component of KnoBo on the final performance, including bottleneck size, concept grounding function, and parameter prior. Additional analyses are available in Appendix C.
Figure 4: Ablation of bottleneck sizes on X-ray datasets. The x-axis is the number of randomly selected concepts (KnoBo) or visual features (Linear Probe).
Knowledge Sources. Besides the empirical results on confounded and unconfounded datasets, we measure the diversity of bottleneck $C$ as $\text{{Diversity}}(C)=\frac{1}{|C|^{2}-|C|}\sum_{c_{i}\in C}\sum^{i\neq j}_{c_{j}\in C}\left(1-\text{sim}(c_{i},c_{j})\right)$, where the $\text{sim}(\cdot)$ is the cosine similarity of concept features encoded by sentence transformer [65]. The Diversity computes the distance between each concept and every other concept in the bottleneck. Table 3 compares different knowledge sources. The retrieval-augmented bottlenecks perform better than those generated by prompting, especially for skin lesions, where more specific knowledge is required because prompting lacks diversity. Across both modalities, PubMed is the best overall, performing better for the X-ray modality than other knowledge sources and among the best for skin lesion modalities. In evaluations by two medical students, information from all knowledge sources is rated as highly relevant and groundable (see Appendix C.4). Moreover, shown in Table 12, our retrieval-augmented concepts are attributable, which allows doctors to verify the source of knowledge.
Bottleneck Size. Figure 4 compares KnoBo and linear probes while varying the number of concepts/features. KnoBo consistently outperforms linear probes across all metrics when given the same quota of features, and KnoBo can obtain good performance with fewer features. This demonstrates that interpretable concept scores have more effective priors than black-box visual features.
@@
Ablations. Table 4 summarizes experiments ablating major components of our approach. Row 2 shows the performance of using dot-products from prompted CLIP models as concepts, which markedly reduces performance. This shows the importance of knowledge grounding in ensuring KnoBo’s effectiveness. However, this step can be simplified as more advanced medical foundation models are available. Row 3 shows performance omitting the parameter prior. It is an important mechanism for constraining the final phase of learning, resulting in consistent OOD improvements.
@@ 7 Conclusion and Limitation
In this paper, we analyze domain-shift problems in medical image analysis and identify a missing medical deep image prior as a main contributor to poor performance. To address this, we introduce knowledge-enhanced bottlenecks (KnoBo) to integrate knowledge priors from medical documents. Across two medical image modalities under various domain shifts, KnoBo significantly improves robustness. KnoBo assumes the availability of medical multimodal datasets, limiting applications to rare conditions. While our work improves robustness, medical experts do not fail in these ways, and they should be used in conjunction with models.
@@ Appendix A Dataset
Table 5 and 6 show the detailed statistics for all 20 datasets evaluated in this paper, where we list the number of classes with class names, the number of images for training, validation, and testing. For confounded datasets, here we explain some details about how we define the confounding factors:
@@ B.3 KnoBo Details
Concept Grounding. We use a language model to annotate 2,000 clinical reports for each concept from the pretraining corpus. To efficiently label reports and achieve a balance of positive and negative examples, we retrieve the top 1,000 reports showing high textual similarity (measured by Sentence Transformer [65]) to the concept as its potential positive examples and randomly sample another 1,000 for potential negatives. We use Flan-T5-XXL [10] as the underlying large language model. Specifically, the annotation task is treated as a next token prediction, similar to the approach proposed by McInerney et al. [54], where we compare the probabilities of the next token being Yes or No to determine if the report contains the concept. Figure 8 shows no big difference in final classification performance when using Flan-T5 versus GPT-4 for annotating concepts on reports.
@@ C.3 Ablate Bottleneck Size for Skin Lesion
Knowledge Source Relevance Groundability
X-ray Skin X-ray Skin
Prompt 3.83 3.93 3.03 3.00
Textbooks 3.70 3.80 2.90 3.27
Wikipedia 3.80 3.67 2.83 3.33
StatPearls 3.87 3.80 2.70 2.97
PubMed 3.70 3.83 2.77 3.20
C.4 Human Evaluation on Bottlenecks
Two medical students evaluated the quality of bottlenecks using two metrics: (1) Relevance measures the concept’s relevance to diagnosing diseases on a scale from 1 (not at all relevant) to 4 (mostly relevant), and (2) Groundability assesses the verifiability of the concept from the image on a scale from 1 to 4. We evaluated 30 randomly sampled concepts from each bottleneck. Table 11 presents these metrics for bottlenecks constructed from five different knowledge sources. While all bottlenecks show good relevance, groundability scores are lower, reflecting the challenge of deriving visual concepts from text-only data.
@@ Appendix D Qualitative Examples
Figure 7 compares the CLIP dot product and our concept grounding functions for retrieving images based on text queries. Our method demonstrates superior recall of correct examples compared to CLIP alignments. Figure 8 showcases the top concepts selected based on the linear weights learned by KnoBo, which are most correlated with the corresponding disease class. These examples highlight that the top concepts utilized by KnoBo are essential features doctors use for diagnosing targeted diseases. Table 12 displays examples of concepts generated by our retrieval-augmented approach. Each concept can be traced back to its source document to verify its accuracy. Figure 9 presents examples of GPT-4V-annotated clinical reports for skin lesion images. Notably, the second example demonstrates GPT-4V’s capability to assess lesion size using the scale provided in the image.
@@
Bottleneck Concept Query Reference Document

### v2

@@ Abstract
While deep networks have achieved broad success in analyzing natural images, when applied to medical scans, they often fail in unexpected situations. We investigate this challenge and focus on model sensitivity to domain shifts, such as data sampled from different hospitals or data confounded by demographic variables such as sex, race, etc, in the context of chest X-rays and skin lesion images. A key finding we show empirically is that existing visual backbones lack an appropriate prior from the architecture for reliable generalization in these settings. Taking inspiration from medical training, we propose giving deep networks a prior grounded in explicit medical knowledge communicated in natural language. To this end, we introduce Knowledge-enhanced Bottlenecks (KnoBo), a class of concept bottleneck models that incorporates knowledge priors that constrain it to reason with clinically relevant factors found in medical textbooks or PubMed. KnoBo uses retrieval-augmented language models to design an appropriate concept space and an automatic training procedure for recognizing the concept. We evaluate different resources of knowledge and recognition architectures on a broad range of domain shifts across 20 datasets. In our comprehensive evaluation with two imaging modalities, KnoBo outperforms fine-tuned models on confounded datasets by 32.4 % on average. Finally, evaluations reveal that PubMed is a promising resource for making medical models less sensitive to domain shift, outperforming other resources on both diversity of information and final prediction performance.
@@ 2 Related Work
Knowledge Rich Multimodal Reasoning. Knowledge plays an important role in clinical diagnosis [6]. Some multimodal tasks [82, 53, 74] require models to use explicit outside knowledge to make correct predictions. Previous methods [52, 50, 31] retrieve documents for each example from the external knowledge base as context for models to generate the answer. Our work focuses on leveraging knowledge in medical image classification. Retrieval-Augmented Generation [48, 23] has been shown to be beneficial for knowledge-intensive tasks [38], including biomedicine [20, 84]. The retrieved medical documents are either used as context during inference [88] or data for pretraining [94]. In contrast, we treat documents as background knowledge for large language models to build concept bottlenecks. Instead of retrieving documents for every input, we build a global knowledge prior from a medical document corpus, which is shared across all examples.
@@ 4 Knowledge-enhanced Bottlenecks
In this section, we present Knowledge-enhanced Bottlenecks (KnoBo), a class of CBMs that incorporate knowledge priors that address the failures we identified in Section 3. Figure 3 presents an overview of our method. From left to right, we optimize three terms: (1) Structure Prior (Sec 4.2) induces bottleneck structures from medical corpus to incorporate human knowledge as concepts, (2) Bottleneck Predictor (Sec 4.3) projects input image onto bottleneck concepts and then feed concept predictions into the linear layer for label prediction, and (3) Parameter Prior (Sec 4.4) aligns the learned parameters with known associative information to further enhance priors.
@@ 4.1 Problem Formulation
$\log P\left(y,C,\theta|I\right)=\underbrace{\log P\left(C\right)}_{\text{structure prior}}+\underbrace{\log P\left(\theta\right)}_{\text{parameter prior}}+\underbrace{\log P\left(y|I,C,\theta\right)}_{\text{bottleneck predictor}}$ (1)
@@
Algorithm 1 Retrieve Augmented Iterative Concept Bottleneck Generation
@@ 4.2 Structural Prior
Given a background corpus $\mathcal{B}$ that spans various documents, we aim to identify a bottleneck structure containing concepts beneficial for classifying labels $y\in\mathcal{Y}$. As outlined in Algorithm 1, we use the class names $\mathcal{Y}$ as initial queries to retrieve relevant documents $\mathcal{B^{\prime}}\subset\mathcal{B}$. Large Language Models (LLMs) are then prompted to generate concepts using these retrieved documents as context: $C^{\prime}=\text{LLM}(\mathcal{B^{\prime}})$. These newly generated concepts are added to the bottleneck and used as new queries to retrieve additional documents. This iterative process continues to expand the bottleneck until a predetermined number of concepts $N_{C}$ is reached. Such concept structures have high likelihood under the language model, conditioned on the background corpus, and the language model probability serves as an implicit prior.
@@ 5.1 Datasets
We evaluate two groups of datasets for each modality: (1) the confounded datasets, which aim to assess the robustness of models by creating splits with spurious correlations; (2) the unconfounded datasets are randomly split to measure the models’ performance in natural settings.
Confounded Datasets. As illustrated on the left of Figure 1, we formulate the confounded datasets as binary classification tasks, where each class is confounded with one factor. The confounding combinations are reversed for in-domain (train and validation), and out-of-domain (test) splits.
@@ 5.3 Implementation Details
KnoBo Details. We select GPT-4 (gpt-4-0613) as the underlying LLM for retrieval-augmented concept generation (Sec 4.2). For training concept grounding functions (Sec 4.3), we opt for Flan-T5-XXL [10] to annotate clinical reports for each concept, considering cost-efficiency. Unless otherwise specified, KnoBo uses bottlenecks constructed from PubMed, each with 150 concepts. Figure 5 shows our prompt, and during concept generation, we apply several heuristic filters (Appendix  B.3).
@@ 6.1 Main Results
KnoBo is more robust to domain shifts. Table 1 shows the results on 10 confounded datasets of X-ray and skin lesions. Black-box models excel at in-domain (ID) data but drop significantly on out-of-domain (OOD) data, especially in datasets confounded by hospitals/resources (NIH-CheXpert and ISIC-hospital), which can be common when collecting medical datasets [12, 81]. KnoBo outperforms baselines in OOD and domain-average accuracy by large margins, ranking top-1 in eight datasets and second-best in the other two. End-to-end models (ViT-L/14, DenseNet) exhibit larger domain gaps than linear probes, as they have more parameters to optimize performance on in-domain data and capture spurious correlations. Shaping the visual representations with knowledge (LSL) improves robustness but underperforms KnoBo, with lower ID, OOD, and average performance across most datasets. PCBM-h combines interpretable and black-box predictions but exhibits behaviors similar to black-box models with severe drops across domains. Unlike KnoBo, which uses medical documents to create one global bottleneck for each modality, LaBo builds a bottleneck for each dataset using the in-domain data, which can be biased and affected by confounding factors and performs more poorly. In summary, KnoBo mitigates the catastrophic failures in domain shifts encountered by black-box and is more robust against various confounding factors across modalities.
@@ 6.2 Analysis
In this section, we compare the bottlenecks constructed from different knowledge sources. We evaluate the impact of each KnoBo component on the final performance, including bottleneck size, concept grounding function, and parameter prior. Additional analyses are available in Appendix C.
Knowledge Sources. Besides the empirical results on confounded and unconfounded datasets, we measure the diversity of bottleneck $C$ as $\text{{Diversity}}(C)=\frac{1}{|C|^{2}-|C|}\sum_{c_{i}\in C}\sum^{i\neq j}_{c_{j}\in C}\left(1-\text{sim}(c_{i},c_{j})\right)$, where the $\text{sim}(\cdot)$ is the cosine similarity of concept features encoded by sentence transformer [65]. The Diversity computes the distance between each concept and every other concept in the bottleneck. Table 3 compares different knowledge sources. The retrieval-augmented bottlenecks perform better than those generated by prompting, especially for skin lesions, where more specific knowledge is required because prompting lacks diversity. Across both modalities, PubMed is the best overall, performing better for the X-ray modality than other knowledge sources and among the best for skin lesion modalities. Moreover, shown in Table 6, our retrieval-augmented concepts are attributable, which allows doctors to verify the source of knowledge.
Knowledge Source Relevance Groundability
X-ray Skin X-ray Skin
Prompt 3.83 3.93 3.03 3.00
Textbooks 3.70 3.80 2.90 3.27
Wikipedia 3.80 3.67 2.83 3.33
StatPearls 3.87 3.80 2.70 2.97
PubMed 3.70 3.83 2.77 3.20
Human Evaluation on Bottlenecks. In evaluations by two medical students, information from all knowledge sources is rated as highly relevant and groundable. Two medical students evaluated the quality of bottlenecks using two metrics: (1) Relevance measures the concept’s relevance to diagnosing diseases on a scale from 1 (not at all relevant) to 4 (mostly relevant), and (2) Groundability assesses the verifiability of the concept from the image on a scale from 1 to 4. We evaluated 30 randomly sampled concepts from each bottleneck. Table 4 presents these metrics for bottlenecks constructed from five different knowledge sources. While all bottlenecks show good relevance, groundability scores are lower, reflecting the challenge of deriving visual concepts from text-only data.
Figure 4: Ablation of bottleneck sizes on X-ray datasets. The x-axis is the number of randomly selected concepts (KnoBo) or visual features (Linear Probe). $+$prior means adding parameter prior.
Bottleneck Size. Figure 4 compares KnoBo and linear probes while varying the number of concepts/features. KnoBo consistently outperforms linear probes across all metrics when given the same quota of features, and KnoBo can obtain better performance with fewer features. This indicates that interpretable concept scores have more effective priors than black-box visual features.
@@
Ablations. Table 5 summarizes experiments ablating major components of our approach. Row 2 shows the performance of using dot-products from prompted CLIP models as concepts, which markedly reduces performance. This shows the importance of knowledge grounding in ensuring KnoBo’s effectiveness. However, this step can be simplified as more advanced medical foundation models are available. Row 3 shows performance omitting the parameter prior. It is an important mechanism for constraining the final learning phase, resulting in consistent OOD improvements. This is also reflected in Figure 4, where with the parameter prior, KnoBo performs better on OOD splits while decreasing on in-domain as the parameter prior constrains the model to rely on data.
Bottleneck Concept Query Reference Document
@@ 7 Conclusion and Limitation
In this paper, we analyze domain-shift problems in medical image analysis and identify a missing medical deep image prior as a main contributor to poor performance. To address this, we introduce knowledge-enhanced bottlenecks (KnoBo) to integrate knowledge priors from medical documents. Across two medical image modalities under various domain shifts, KnoBo significantly improves robustness over black-box baselines.
Limitation. KnoBo assumes the availability of medical multimodal datasets, limiting applications to rare conditions. While our work improves robustness, medical experts do not fail in these ways, and they should be used in conjunction with models. Our preliminary work with PubMed suggests it is a valuable resource for developing medical models, and future research can explore how to utilize such fruitful knowledge resources more effectively.
Acknowledgement
This study was conducted at the University of Pennsylvania School of Engineering and Applied Science and supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005, and gifts from the UPenn ASSET center and Ai2. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes, notwithstanding any copyright annotation therein. Michael S. Yao was supported by the National Institutes of Health (F30 MD020264). James C. Gee was also supported by the NIH (R01 EB031722).
@@ Appendix A Dataset
Table 7 and 8 show the detailed statistics for all 20 datasets evaluated in this paper, where we list the number of classes with class names, the number of images for training, validation, and testing. For confounded datasets, here we explain some details about how we define the confounding factors:
@@ B.3 KnoBo Details
Concept Grounding. We use a language model to annotate 2,000 clinical reports for each concept from the pretraining corpus. To efficiently label reports and achieve a balance of positive and negative examples, we retrieve the top 1,000 reports showing high textual similarity (measured by Sentence Transformer [65]) to the concept as its potential positive examples and randomly sample another 1,000 for potential negatives. We use Flan-T5-XXL [10] as the underlying large language model. Specifically, the annotation task is treated as a next token prediction, similar to the approach proposed by McInerney et al. [54], where we compare the probabilities of the next token being Yes or No to determine if the report contains the concept. Table 10 shows no big difference in final classification performance when using Flan-T5 versus GPT-4 for annotating concepts on reports.
@@ C.3 Ablate Bottleneck Size for Skin Lesion
Table 13: Ablate the number of clinical reports for training each concept grounding function.
Reports / Concept Confd Unconfd Overall
100 72.5 71.7 72.1
250 73.7 72.1 72.9
500 73.1 73.0 73.1
1000 73.1 72.8 73.0
1500 73.5 73.2 73.4
2000 74.3 73.1 73.7
C.4 Data Efficiency for Concept Grounding
The concept grounding module of KnoBo requires image-report pairs from a multimodal medical dataset. By default, we select 2,000 pairs for training the grounding function for each concept. However, this number can be largely reduced. Table 13 shows KnoBo’s performance doesn’t decrease too much with fewer training examples. With only 100 reports, KnoBo can obtain 72.1 overall accuracy, compared to 73.7, of which 2,000 reports are used. It highlights that KnoBo achieves much of its performance with a smaller number of examples, suggesting that it is not data-hungry and has the potential to be applied to rare modalities.
@@ Appendix D Qualitative Examples
Figure 7 compares the CLIP dot product and our concept grounding functions for retrieving images based on text queries. Our method demonstrates superior recall of correct examples compared to CLIP alignments. Figure 8 showcases the top concepts selected based on the linear weights learned by KnoBo, which are most correlated with the corresponding disease class. These examples highlight that the top concepts utilized by KnoBo are essential features doctors use for diagnosing targeted diseases. Table 14 displays examples of failed concepts identified by medical students during human evaluation. Each concept can be traced back to its source document to verify its accuracy. Figure 9 presents examples of GPT-4V-annotated clinical reports for skin lesion images. Notably, the second example demonstrates GPT-4V’s capability to assess lesion size using the scale provided in the image.
@@
Modality Irrelevant Concepts Ungroundable Concepts Repetitive Concepts
X-ray $\bullet$ Are there any subcutaneous emphysema? $\bullet$ Are there changes in appearance that develop slowly? $\bullet$ Is there evidence of ground-glass opacity?
$\bullet$ Can you spot elevated hems in the diaphragm? $\bullet$ Can worsening consolidation be observed? $\bullet$ Is there ground glass opacity in both lung fields?
Skin $\bullet$ Is the lesion located on the surface of extremities? $\bullet$ Is the skin around the lesion rough without pain? $\bullet$ Does the lesion have regular margins?
$\bullet$ Does it have scales near the lesion? $\bullet$ Is the lesion larger than 6mm in diameter? $\bullet$ Does the lesion possess an irregular shape?
$\bullet$ Are skin lesions involving the face and scalp? $\bullet$ Is there any change in the size of the lesion over time? $\bullet$ Does the lesion have irregular borders?
Table 14: We let medical students annotate 300 concepts and categorize the failure cases into three categories: (1) Irrelevant: concepts are not essential to diagnosis; (2) Ungroundable: concepts cannot be grounded on only one image and (3) Repetitive: the concepts are repeated in the bottleneck.